# SyncGuard: Robust Audio Watermarking Capable of Countering Desynchronization Attacks

### v1

@@ 2.1 Traditional Audio Watermarking
Traditional audio watermarking primarily embeds watermark information in the time domain and frequency domain. Time-domain methods embed watermark by directly altering the values of the audio signal. For example, Xiang et al. [23] propose a novel dual-channel time-spread echo method for audio watermarking. Compared to time-domain methods, frequency-domain methods can further enhance robustness and achieve better perceptual quality. These methods embed watermark information by slightly adjusting the frequency coefficients using Fourier transforms, such as Discrete Cosine Transform (DCT) [8, 2, 17] and Discrete Wavelet Transform (DWT) [1, 5].
Watermark localization and resistance to desynchronization attacks remain challenging issues. To address these challenges, Liu et al. [16] embed synchronization codes into the residuals of frequency domain coefficients and employ exhaustive searches with sliding windows of different sizes to achieve synchronization. Liu et al. [16] introduces frequency-domain coefficients logarithmic mean (FDLM) features. It embeds synchronization codes at the beginning of each audio frame to improve alignment robustness. However, the decoding process relies on fixed-length segmentation, making it less effective when audio length is altered by attacks such as time-scaling or cropping. For stereo audio signals, Zong et al. [27] modify the Pearson Correlation Coefficient (PCC) between the DCT coefficients of left and right channel signals. Zhao et al. [25] introduced a novel feature termed Frequency Singular Value Coefficient (FSVC) extracted from DCT domain. Zhao et al. [26] utilize Segmental Singular Value Summation (SSVS) and Segmental Singular Value Difference (SSVD) features as information carriers and design an adaptive method to select embedding parameters. Although these methods can resist desynchronization attacks, the manual design approach limits their flexibility, posing challenges in balancing robustness, audio quality, and embedding capacity.
@@ 3.1 Watermark Embedding
s,p=STFT(a).s,p={STFT}(a). (1)
We use a linear layer $FC$ to expand the $n$ bits of the watermark into a vector $f_{w}$, the same size as one frame of the spectrogram $s$.
@@
Frame-wise broadcast. Previous methods[13, 9] directly extend the watermark vector through linear layers to match the shape of the waveform or features, sacrificing temporal flexibility and leading to a non-uniform distribution of watermark information along the temporal axis. In our approach, we perform a frame-wise broadcast mechanism of the watermark feature $f_{w}$ along the time domain until its temporal dimension matches that of $s$, resulting in $f_{wb}$.
@@
$a_{w}={ISTFT}\left(s_{w},p\right).$ (6)
@@ 3.2 Watermark Extraction
Initially, we utilize Eq.(1) to perform the STFT on $a_{w}$ to derive the phase information $p_{w}$ and spectrogram $s_{w}$.
$s_{w},p_{w}={STFT}(a_{w}).$ (7)
@@ 3.4.1 Time Scale Modification
$s^{\prime}=Timewarp(s,rate),$ (12)
where rate denotes the time-stretch ratio. The function $Timewarp(\cdot)$ denotes a differentiable interpolation function. Meanwhile, for each frequency component $f$, we predict the phase information $p^{\prime}_{f}(t)$ at time $t$.
$p^{\prime}_{f}(t)=p_{f}(t)+\Delta p_{f}(t).$ (13)
@@
$\Delta p_{f}(t)={unwrap}(p_{f}(t+1)-p_{f}(t))\cdot r.$ (14)
The function ${unwrap}(\cdot)$ is used to handle the periodicity of the phase, ensuring that the phase difference remains within the range of $\pi$ to $\pi$. Here, $r$ represents the time scale modification factor. Subsequently, the warped audio signal $a_{time}$ is obtained by applying the $ISTFT$ described in Eq. (6) to the distorted phase and spectrogram.
$a_{time}=ISTFT(s^{\prime},p^{\prime}).$ (15)
@@ 3.4.2 Pitch Scaling
Specifically, the process begins with the application of time scale modification, where the scaling factor is calculated based on the number of semitones the pitch is shifted.
$\begin{gathered}{rate}=2^{\frac{x}{12}},\\ a_{s}=Timewarp(a_{w},rate).\end{gathered}$ (16)
@@ 3.5 Training Strategy
$\mathcal{L}=\lambda_{e}\cdot\mathcal{L}_{e}+\lambda_{adv}\cdot\mathcal{L}_{adv}+\lambda_{w}\cdot\mathcal{L}_{w}.$ (18)
@@ 4.1.4 Comparative Methods
We compared the proposed SyncGuard with state-of-the-art methods, including FSVC [25], FDLM [16], DeAR [13] and DRAW [9]. It is important to note that the first two methods do not require any training. For DeAR, due to its lack of convergence at high embedding rates, we adhered to its settings and embedded 100 bits. For DRAW, we followed its settings, embedding at an effective capacity of 32.73 bps, while the embedding capacity for other methods remained at 32 bps.
@@ 4.2 Experimental Results
Figure 5: Visualization of a watermarked audio spectrogram (log-frequency scale, amplitude in dB).
@@ 4.2.2 Robustness Against Common Signal Processing
Table 2: Robustness against common signal processing. The symbol * indicates the attack (or corresponding parameter) was not included during training.
@@
Table 3: Robustness Against Desynchronization Attacks. The symbol * indicates the attack (or corresponding parameter) was not included during training.
@@ 4.2.3 Robustness Against Desynchronization Attacks
We employ the following desynchronization attacks to assess the robustness of our proposed method and the methods chosen for comparison:
(1) Jittering attack: The watermarked audio are randomly deleted one sample from every 100 consecutive samples.
(2) Cropping:The watermarked audio was cropped at random positions, with 10% and 20% of the audio removed.
(3) TSM: Time scale the watermarked audio with attack factors of $0.8,0.9,1.1$, and $1.2$.
(4) PS: Pitch scale the watermarked audio with attack factors of $0.9$ and $1.1$.
@@
To further investigate SyncGuard’s robustness against cropping, we conducted tests by cropping the watermarked audio at three different positions: the beginning, middle, and end, and assessed the ACC under varying cropping ratios. As depicted in Figure 6, even with 85% of the audio cropped, our method maintains a 100% extraction accuracy. This superior performance is attributed to our frame-wise broadcast embedding strategy, which ensures that the embedded watermark is uniformly and comprehensively distributed across all parts of the audio, thereby providing strong resilience against cropping.
@@ 4.2.4 Flexibility
Figure 7: Performance at different time offsets from the watermark insertion position. ++ indicates a rightward shift in time, while - indicates a leftward shift.

### v2

@@ 2.1 Traditional Audio Watermarking
Traditional audio watermarking primarily embeds watermark information in the time domain and frequency domain. Time-domain methods embed watermark by directly altering the values of the audio signal. For example, Xiang et al. [22] propose a novel dual-channel time-spread echo method for audio watermarking. Compared to time-domain methods, frequency-domain methods can further enhance robustness and achieve better perceptual quality. These methods embed watermark information by slightly adjusting the frequency coefficients using Fourier transforms, such as Discrete Cosine Transform (DCT) [6, 15] and Discrete Wavelet Transform  [3].
Watermark localization and resistance to desynchronization attacks remain challenging issues. To address these challenges, Liu et al. [14] introduces frequency-domain coefficients logarithmic mean (FDLM) features. It embeds synchronization codes at the beginning of each audio frame to improve alignment robustness. However, the decoding process relies on fixed-length segmentation, making it less effective when audio length is altered by attacks such as time-scaling or cropping. For stereo audio signals, Zong et al. [26] modify the Pearson Correlation Coefficient (PCC) between the DCT coefficients of left and right channel signals. Zhao et al. [24] introduced a novel feature termed Frequency Singular Value Coefficient (FSVC) extracted from DCT domain. Zhao et al. [25] utilize Segmental Singular Value Summation (SSVS) and Segmental Singular Value Difference (SSVD) features as information carriers and design an adaptive method to select embedding parameters. Wu et al. [21] propose LIPAS, leveraging local invariant points and adaptive strength for robust, imperceptible watermarking. Although these methods can resist desynchronization attacks, the manual design approach limits their flexibility, posing challenges in balancing robustness, audio quality, and embedding capacity.
@@ 3.1 Watermark Embedding
$s,p={STFT}(a)\text{.}$ (1)
We use a linear layer $FC$ to expand the $n$ bits of the watermark into a vector $f_{w}$, the same size as one frame of the spectrogram $s$:
@@
Frame-wise broadcast. Previous methods[11, 7] directly extend the watermark vector through linear layers to match the shape of the waveform or features, sacrificing temporal flexibility and leading to a non-uniform distribution of watermark information along the temporal axis. In our approach, we perform a frame-wise broadcast mechanism of the watermark feature $f_{w}$ along the time domain until its temporal dimension matches that of $s$, resulting in $f_{wb}$:
@@
$a_{w}={ISTFT}\left(s_{w},p\right)\text{.}$ (6)
@@ 3.2 Watermark Extraction
Initially, we utilize Eq.(1) to perform the STFT on $a_{w}$ to derive the phase information $p_{w}$ and spectrogram $s_{w}$:
$s_{w},p_{w}={STFT}(a_{w})\text{.}$ (7)
@@ 3.4.1 Time Scale Modification
$s^{\prime}=Timewarp(s,rate)\text{,}$ (12)
where rate denotes the time-stretch ratio. The function $Timewarp(\cdot)$ denotes a differentiable interpolation function. Meanwhile, for each frequency component $f$, we predict the phase information $p^{\prime}_{f}(t)$ at time $t$:
$p^{\prime}_{f}(t)=p_{f}(t)+\Delta p_{f}(t)\text{.}$ (13)
@@
$\Delta p_{f}(t)={unwrap}(p_{f}(t+1)-p_{f}(t))\cdot r\text{.}$ (14)
The function ${unwrap}(\cdot)$ is used to handle the periodicity of the phase, ensuring that the phase difference remains within the range of $\pi$ to $\pi$. Here, $r$ represents the time scale modification factor. Subsequently, the warped audio signal $a_{time}$ is obtained by applying the $ISTFT$ described in Eq. (6) to the distorted phase and spectrogram:
$a_{time}=ISTFT(s^{\prime},p^{\prime})\text{.}$ (15)
@@ 3.4.2 Pitch Scaling
Specifically, the process begins with the application of time scale modification, where the scaling factor is calculated based on the number of semitones the pitch is shifted:
$\begin{gathered}{rate}=2^{\frac{x}{12}},\\ a_{s}=Timewarp(a_{w},rate)\text{.}\end{gathered}$ (16)
@@ 3.5 Training Strategy
$\mathcal{L}=\lambda_{e}\cdot\mathcal{L}_{e}+\lambda_{adv}\cdot\mathcal{L}_{adv}+\lambda_{w}\cdot\mathcal{L}_{w}\text{.}$ (18)
@@ 4.1.4 Comparative Methods
We compared the SyncGuard with state-of-the-art methods, including FSVC [24], FDLM [14], DeAR [11] and DRAW [7]. It is important to note that the first two methods do not require any training. For DeAR, due to its lack of convergence at high embedding rates, we adhered to its settings and embedded 100 bits. For DRAW, we followed its settings, embedding at an effective capacity of 32.73 bps, while the embedding capacity for other methods remained at 32 bps.
@@ 4.2 Experimental Results
Figure 5: Visualization of a watermarked audio spectrogram.
@@ 4.2.2 Robustness Against Common Signal Processing
Table 2: Robustness against common signal processing. * indicates the attack (or corresponding parameter) was not included during training.
@@
Table 3: Robustness against desynchronization attacks. * indicates the attack (or corresponding parameter) was not included during training.
@@ 4.2.3 Robustness Against Desynchronization Attacks
We employ the following desynchronization attacks to assess the robustness of our proposed method and the methods chosen for comparison: (1) Jittering attack: The watermarked audio are randomly deleted one sample from every 100 consecutive samples. (2) Cropping:The watermarked audio was cropped at random positions, with 10% and 20% of the audio removed. (3) TSM: Time scale the watermarked audio with attack factors of $0.8,0.9,1.1$, and $1.2$. (4) PS: Pitch scale the watermarked audio with attack factors of $0.9$ and $1.1$.
@@
To further investigate SyncGuard’s robustness against cropping, we conducted tests by cropping the watermarked audio at three different positions: the beginning, middle, and end, and assessed the ACC under varying cropping ratios. As depicted in Figure 6, even with 85% of the audio cropped, our method maintains a 100% accuracy. This superior performance is attributed to our frame-wise broadcast embedding strategy, which ensures that the embedded watermark is uniformly and comprehensively distributed across all parts of the audio, thereby providing strong resilience against cropping.
@@ 4.2.4 Flexibility
Figure 7: Performance at different time offsets from the watermark insertion position. Positive values on the time axis indicate a rightward shift, while negative values indicate a leftward shift.